### v1

# IntelliProof: An Argumentation Network-based Conversational Helper for Organized Reflection

Abstract
We present IntelliProof, an interactive system for analyzing argumentative essays through LLMs. IntelliProof structures an essay as an argumentation graph, where claims are represented as nodes, supporting evidence is attached as node properties, and edges encode supporting or attacking relations. Unlike existing automated essay scoring systems, IntelliProof emphasizes the user experience: each relation is initially classified and scored by an LLM, then visualized for enhanced understanding. The system provides justifications for classifications and produces quantitative measures for essay coherence. It enables rapid exploration of argumentative quality while retaining human oversight. In addition, IntelliProof provides a set of tools for a better understanding of an argumentative essay and its corresponding graph in natural language, bridging the gap between the structural semantics of argumentative essays and the user’s understanding of a given text. A live demo and the system are available here to try: https://intelliproof.vercel.app
Introduction & Related Work
The rise of Large Language Models (LLMs) has drastically accelerated research in computational argumentation and automated writing support. Argumentative writing is uniquely challenging, requiring a balance of claims, supporting evidence, and counterarguments within a coherent, persuasive structure. Traditional analysis methods, from rule-based systems to neural encoders, frequently struggle to capture the nuanced interrelations between claims and evidence (Elaraby and Litman 2022).
Figure 1: IntelliProof user interface overview using an example graph on the effect of green space on urban environments.
We introduce IntelliProof, an LLM-powered tool that analyzes arguments by modeling them as graphs (Saveleva et al. 2021). In this model, claims are represented as nodes, with their strength quantified by evidence encoded as node properties. Weighted edges denote support or attack relations between claims. An LLM is used to score, classify, and justify these relations, while allowing human overrides for transparency and control. The dynamic identification and visualization of these relationships are shown in Figure 1.
By transforming essays into structured argumentation graphs, IntelliProof aims to make argumentative reasoning more interpretable, providing writers and educators with insights into essay coherence and persuasiveness. This approach contributes to the discussions on how to integrate LLMs into workflows that demand interpretability, reliability, and pedagogical value simultaneously.
LLMs have shifted argument mining methods from encoder-based architectures to prompting and fine-tuning strategies (Cabessa, Hernault, and Mushtaq 2024; Favero et al. 2025). However, annotation bottlenecks and evaluation challenges remain (Schaefer 2025). Recent work also explores interactive systems that combine generative models with human input for constructing argument graphs (Lenz and Bergmann 2025). IntelliProof extends this work by integrating graph-based structuring directly into analysis while grounding scoring of arguments in quantifiable, mathematical metrics.
Educational applications increasingly use LLMs for essay scoring and feedback (Kim and Jo 2024; Chu et al. 2025). Although many approaches optimize predictive accuracy, few address the interpretability of argumentative quality. Surveys of persuasive applications highlight both the promise and ethical risks of LLM-driven reasoning systems (Rogiers et al. 2024). By grounding essay feedback in explicit argument graphs, IntelliProof contributes to more interpretable educational tools, which will lead to safer AI systems deployed in educational settings.
Intelliproof Overview
Intelliproof’s functionality spans argument creation, scoring, classification, and generation techniques. Each of the features elaborated on below is integrated within our GUI front-end.
Graph Visualization
IntelliProof is designed to structurally visualize argumentative essays while providing an LLM-powered (GPT-4o for the instance of the demo given its performance (Shahriar et al. 2024)) toolset for the analysis of the claims. As such, users can input claims, classify them (into Fact, Policy, or Value), and establish connections between the claims via the main GUI of the tool.
LLM Document Analysis
To establish claims, users upload supporting documents as evidence in PDF or image format. A dedicated LLM instance then processes these files, suggesting relevant text or image extracts for a specific claim. The user attaches the suggested evidence to the claim via a drag-and-drop interface, which in turn prompts the LLM to assess the claim’s strength by analyzing all attached evidence. Any number of supporting or negating evidence pieces can be associated with a single claim.
Claim Credibility Score
To assess overall claim strength, we combine evidence and edge scores to obtain the claim credibility score $S_{t}$ where $S_{t}=\tanh(\Delta\frac{1}{n}\Sigma_{i=0}^{n}f_{E}(e_{i})+\Sigma_{j=0}^{m}f_{ED}(k_{j})*S_{t-1})$. $f_{E}$ and $f_{ED}$ are calculated based on the LLMs assessment of claim support based on an evidence, and based on an incoming edge, respectively, and $\Delta$ is a tunable hyperparameter. Note that given the weakness of LLMs in directly generating scores (Schroeder and Wood-Doughty 2025; Cui 2025), we first generate a qualitative classification as the LLM’s assessment, and then utilize the Evans coefficient interpretation (Evans 1996) to convert the qualitative assessment to numerical scores.
Report Generation
Another feature of Intelliproof is automatic report generation from the graph implementation of an argument. These reports combine evidence evaluation, edge validation, assumptions analysis, and graph critique into a singular unified report. Our system processes graph structure, evidence quality, relationship strengths, and logical patterns simultaneously and creates an eight section, comprehensive report of the argumentative essay.
AI Copilot Chat Interface
Using our integrated chatbot, natural language queries are parsed, and one may ask questions about the graph. The responding AI is context aware, and users can get insights on arguments’ strengths, weaknesses, and gaps to fill. As arguments are built, the LLM context window is also updated in real-time to contain the new information.
Assumption Generation
Intelliproof analyzes claim relationships to identify three implicit assumptions that would strengthen support between claims. It also finds hidden premises and bridges assumptions needed to make arguments more robust. Each assumption includes an importance rating and a justification for why it strengthens the relationship generated based on a few-shot learning approach (Brown et al. 2020) prepared by an argumentation field expert.
Critique Graph
To identify essay weaknesses, we deploy a state-of-the-art LLM (GPT-4o) to match the overall argument against our comprehensive Argument Patterns Bank. This Bank is a built-in YAML database, developed by an argumentation expert, containing patterns for logical fallacies, good arguments, and absurd reasoning. This process allows us to specifically identify issues like circular reasoning, straw man arguments, and false causes.
System Implementation
IntelliProof’s architecture consists of three core components. The frontend is built with Vite and React.js to create a dynamic user interface that handles all back-end API and database calls. The backend uses a Python server with FastAPI for handling requests and SupaBase (PostgreSQL) for managing user data such as profiles, evidence files, and graphs. For the large language model, we utilize GPT-4o via OpenAI’s Python library, chosen for its balance of performance, cost, and availability. The design is modular, allowing GPT-4o to be easily substituted with other locally or remotely deployed LLMs. [^1] We publicly release the source code for IntelliProof at https://github.com/collective-intelligence-lab/intelliproof
Conclusion
Intelliproof is an interactive LLM platform that creates argument graphs based on provided evidence. This system is designed to be helpful in devising strong arguments, filling gaps in arguments, and utilizing an LLM to provide a detail-oriented look at an argumentative essay. While this system may be expanded further in the future, at present, we provide a robust, functional system that demonstrates the feasibility of Intelliproof as a powerful tool for structured, LLM-driven argumentation.
Acknowledgments
This research was supported in part by Other Transaction award HR00112490378 from the U.S. Defense Advanced Research Projects Agency (DARPA) Friction for Accountability in Conversational Transactions (FACT) program.